### v1

@@ 1 Preliminaries: the algebraic classification
$\mathcal{L}_{02}$ : $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{03}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{04}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{05}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{06}$ :::: $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}$.
$\mathcal{L}_{07}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{08}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{09}$ :::: $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{10}$ :::: $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{11}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{12}$ :::: $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{13}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{14}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{15}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{16}^{\alpha}$ :::: $[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{17}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$ $[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{18}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$
$[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{19}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$
$\mathcal{L}_{20}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{21}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4}.$
$\mathcal{L}_{22}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{23}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{24}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{25}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{26}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{27}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{28}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{29}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{30}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{31}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{32}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{33}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{34}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{35}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{36}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{37}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{38}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{39}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{40}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{41}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{42}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=-e_{4},$
$[e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{43}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$
$[e_{1},e_{2},e_{3}]=-e_{4},$ $[e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{44}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3}+e_{4},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$
$[e_{1},e_{2},e_{3}]=-e_{4},$ $[e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{45}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{46}^{\alpha\neq 0}$ :::: $[e_{1},e_{2}]=e_{3}+\alpha e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{47}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4},$
$[e_{1},e_{3},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{48}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{49}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{50}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{51}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{52}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
@@
Definition 7 (see [12]).
A Bol algebra is a vector space $\mathfrak{B}$ with a bilinear multiplication [,][-,-][ - , - ] and a trilinear multiplication [,,][-,-,-][ - , - , - ] satisfying:
$\begin{array}[]{ll}({\rm B1})&[x,y]+[y,x]=0,\\ ({\rm B2})&[x,y,z]+[y,x,z]=0,\\ ({\rm B3})&[x,y,z]+[y,z,x]+[z,x,y]=0,\\ ({\rm B4})&[u,v,[x,y,z]]-[[u,v,x],y,z]-[x,[u,v,y],z]-[x,y,[u,v,z]]=0,\\ ({\rm B5})&[[x,y,z],t]-[[x,y,t],z]+[z,t,[x,y]]-[x,y,[z,t]]+[[x,y],[z,t]]=0.\\ \end{array}$

### v2

@@ 1 Preliminaries: the algebraic classification
Example 1.4.
The Lie–Yamaguti algebra with $[x,y]=0$ for any $x,y\in\mathcal{L}$ is exactly a Lie triple system, closely related to symmetric spaces, while the Lie–Yamaguti algebra with $[x,y,z]=0$ for any $x,y,z\in\mathcal{L}$ is a Lie algebra.
Theorem 1.5 (see [1]).
Let $\mathcal{L}$ be a $4$-dimensional complex nilpotent Lie–Yamaguti algebra. Then $\mathcal{L}$ is isomorphic to one of the following pairwise non-isomorphic algebras:
$\mathcal{L}_{01}$ :: $e_{1},e_{2}]=e_{3},$ $\mbox{(Lie algebra)}.$
$\mathcal{L}_{02}$ :: $e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{03}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{04}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{05}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{06}$ :: $e_{2},e_{3},e_{2}]=e_{4}, [$$e_{3},e_{1},e_{3}]=e_{4}$.
$\mathcal{L}_{07}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}, [$$e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{08}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}, [$$e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{09}$ :: $e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}, [$$e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{10}$ :: $e_{2},e_{3},e_{1}]=e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}. [$$e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{11}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4},$
$e_{2},e_{3},e_{2}]=e_{4}. [$$e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{12}$ :: $e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4},$
$e_{2},e_{3},e_{2}]=e_{4}. [$$e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{13}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4},$
$e_{2},e_{3},e_{2}]=e_{4}. [$$e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{14}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4},$
$e_{2},e_{3},e_{2}]=e_{4}. [$$e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{15}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4},$
$e_{2},e_{3},e_{2}]=e_{4}, [$$e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{16}^{\alpha}$ :: $e_{2},e_{3},e_{1}]=\alpha e_{4}, [$$e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{17}^{\alpha}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4}, [$$e_{2},e_{3},e_{1}]=\alpha e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{18}^{\alpha}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$
$e_{2},e_{3},e_{1}]=\alpha e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{19}^{\alpha}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3}]=e_{4}, [$$e_{3},e_{1},e_{2}]=e_{4}.$
$e_{2},e_{3},e_{1}]=\alpha e_{4}, [$$e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$
$\mathcal{L}_{20}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{21}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{4}.$
$\mathcal{L}_{22}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{23}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=\alpha e_{4}, [$$e_{1},e_{2},e_{1}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{24}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{25}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{26}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4}, [$$e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{27}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3},e_{2}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{28}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{3},e_{2}]=e_{4}, [$$e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{29}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{3},e_{2}]=e_{4},$
$e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{30}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{31}$ :: $e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{32}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{33}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{3}]=e_{4}, [$$e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{34}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{3}]=e_{4},$
$e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{35}$ :: $e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}, [$$e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{36}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}, [$$e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{37}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}, [$$e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{38}$ :: $e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{39}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{40}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{41}$ :: $e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{42}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{3}]=-e_{4},$
$e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{43}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3},$
$e_{1},e_{2},e_{3}]=-e_{4}, [$$e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{44}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}+e_{4}, [$$e_{1},e_{3}]=\alpha e_{4}, [$$e_{2},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3},$
$e_{1},e_{2},e_{3}]=-e_{4}, [$$e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{45}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{46}^{\alpha\neq 0}$ :: $e_{1},e_{2}]=e_{3}+\alpha e_{4}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{47}^{\alpha}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{3}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{2}]=e_{4},$
$e_{1},e_{3},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{48}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{49}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{2}]=e_{4}, [$$e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{50}$ :: $e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{51}$ :: $e_{1},e_{2}]=e_{4}, [$$e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{52}$ :: $e_{1},e_{2}]=e_{3}, [$$e_{1},e_{2},e_{1}]=e_{3}, [$$e_{1},e_{2},e_{2}]=e_{4}.$
@@
Definition 1.7 (see [12]).
A Bol algebra is a vector space $\mathfrak{B}$ with a bilinear multiplication [,][-,-] and a trilinear multiplication [,,][-,-,-] satisfying: